OpenStudy (anonymous):

Show that the set T ={(w,x,y,z)∈R4 such that y=w and x^2 =z^4} is not the graph of any function of w and x.

OpenStudy (helder_edwin):

well u have the correspondence (i don't know if this is the word in english) \[ \large (w,x)\mapsto(y,z) \] such that \(y=w\) and \(z^4=x^2\).

OpenStudy (helder_edwin):

do u remember the definition of function?

OpenStudy (anonymous):

A function is a rule that assigns a unique element in Rn to Rm. It is one to one.

OpenStudy (helder_edwin):

the last part is something else.

OpenStudy (helder_edwin):

a function assigns a UNIQUE element to EVERY element of its domain.

OpenStudy (anonymous):

Ok I understand that.

OpenStudy (anonymous):

So a certain w and x cannot have two outputs.

OpenStudy (helder_edwin):

what does (w,x)=(1,-1) get assigned to?

OpenStudy (helder_edwin):

yes. precisely.

OpenStudy (anonymous):

(1,-1) is assigned to (1,1)?

OpenStudy (helder_edwin):

just that?

OpenStudy (helder_edwin):

let's see: \[ \large y=w=1 \] right?

OpenStudy (anonymous):

right

OpenStudy (helder_edwin):

BUT \[ \large z^4=x^2=(-1)^2=1\Rightarrow z=\pm\sqrt[4]{1}=\pm1 \]

OpenStudy (helder_edwin):

so \[ \large (1,-1)\mapsto(1,1) \] and \[ \large (1,-1)\mapsto(1,-1) \]

OpenStudy (anonymous):

So does that mean T is not a function?

OpenStudy (helder_edwin):

yes. that's what u were asked to prove.

OpenStudy (anonymous):

T is graph of something, but that something is not a function. right?

OpenStudy (helder_edwin):

yes.

OpenStudy (anonymous):

Thank you SO much!

OpenStudy (helder_edwin):

u r welcome.